OpenStudy (anonymous):

an indoor track is to be designed such that each end is banked semi-circle with a radius of 24m. what should the banking angle be for a person running at a speed of 6.0m/s?

OpenStudy (anonymous):

 Notice this question involves centripetal force. From the diagram, it shows that the centripetal force, \(F_c={F_N} sin \theta\) And since \(F_c=m\frac{v^2}{r}\), where v and r are given, it should be appropriate that you start by solving \(F_N\) Also note that \(F_g cos \theta = F_N\) However I think the question misses the weight of the person.